OpenStudy (alyssajobug):

Solve the equation for the values of x. |4 - 8x| - 5 = 19

jigglypuff314 (jigglypuff314):

to solve for x for: |4 - 8x| - 5 = 19 you want to get the absolute value alone first +5 +5 |4 - 8x| = 24 which makes 4 - 8x = 24 and 4 - 8x = -24 solve for x for both :)